Suppose HTC sells 8000 smartphones in Taiwan per week at a price of $250.Due to competition,HTC reduces the price to $225 and subsequently notices that demand for its smartphone increases to 8640 units in the following week.Calculate the price elasticity of demand for HTC smartphones
1
Expert's answer
2020-12-04T09:31:22-0500

The price elasticity of demand for HTC smartphones is:

"Ed = \frac{8640 - 8000} {225 - 250} ×\frac {225 + 250} {8640 + 8000} = -0.73."